# Unified Thinker: A General Reasoning Modular Core for Image Generation

### v1

@@ 1 Introduction
The rapid evolution of diffusion-based foundation models ho2020denoising; dhariwal2021diffusion; rombach2022high; xu2024diffusion_dpo has driven an unprecedented leap in high-fidelity image synthesis. Advanced proprietary models such as GPT-4o hurst2024gpt4o and Nano Banana nanobanana have recently demonstrated strong reasoning-driven image generation under complex instructions. In contrast, despite steady progress in open-source systems sd3.5; blackforest2024flux; bagel2025; qwenimage-edit; liao2025imagegen-cot, current open-source models still exhibit a clear gap in handling logic-intensive or implicit directives wisebench; risebench; reason2edit; liu2025step1x-edit.
Current attempts to bridge this gap follow two primary approaches. Built-in Reasoning internalizes reasoning into the generator via unified training that couples multimodal understanding with generation bagel2025; show-o2; xiao2024omnigen. However, this tight entanglement reduces modularity and may destabilize training, often degrading the generator’s visual fidelity. In contrast, External Planner-Driven methods use an MLLM to plan for a mostly frozen generator qwenimage-edit; lin2025uniworld; li2025editthinker; reason2edit. While modular, they suffer from a reasoning–execution mismatch: text-space plans are not grounded in the generator’s capabilities, so even correct plans can cause visual failures, and iterative planning further increases compute.
@@
However, this decoupled design still faces additional challenges: as shown in Fig. 1, without proper alignment, a naive Thinker may produce plausible reasoning that the Generator cannot execute. To bridge the reasoning-to-execution gap, we introduce a dedicated data-to-training pipeline to align planning with visual outcomes. We first construct HieraReason-40K, a hierarchical reasoning dataset synthesized with Gemini-3-Pro nanobanana, which pairs complex instructions with structured, executable plans to teach the Thinker the desired planning format and basic logical decomposition. We then adopt a two-stage training strategy: we perform joint supervised fine-tuning on HieraReason-40K to establish initial plan quality, followed by an end-to-end dual-phase reinforcement learning procedure that places the Generator in the loop and optimizes the Thinker using rewards computed from the final image’s constraint satisfaction. This directly grounds the Thinker’s policy in pixel-level feedback, encouraging plans that are not only semantically plausible but also executable under the Generator’s capabilities.
@@ 2.1 Foundational Generative Models
Modern image generation is predominantly anchored in diffusion-based frameworks ho2020denoising; rombach2022high. Recent advances sd3.5; blackforest2024flux; qwenimage-edit build upon Diffusion Transformers peebles2023scalable and flow matching flowmatch to improve fidelity, prompt alignment, and diversity in latent diffusion models. Meanwhile, an emerging direction unifies autoregressive modeling with visual generation in a single framework, giving rise to unified multimodal models bagel2025; janus; show-o2; xiao2024omnigen. For instance, Bagel bagel2025 uses a transformer backbone to jointly model text and image tokens, whereas OmniGen xiao2024omnigen dispenses with external encoders and handles multiple vision tasks through a unified pipeline. In parallel, image editing has evolved from mask-based inpainting zhuang2024task; ju2024brushnet to instruction-guided manipulation brooks2023instructpix2pix; yu2025anyedit. To further enhance instruction following, recent methods li2024smartedit; fu2024guiding; lin2025uniworld; liu2025step1x-edit such as Qwen-Image-Edit qwenimage-edit leverage MLLMs for instruction parsing and planning. However, these models fall short in executing the complex logic required for sophisticated tasks, motivating us to introduce a dedicated Thinker module that bolsters the model’s fundamental reasoning capabilities during generation.
@@ 2.2 Reasoning for Image Generation
Recent research has moved beyond the one-shot mapping paradigm by explicitly incorporating reasoning into the image generation process. One line of work t2ir1; wang2025mint; liao2025imagegen-cot; unicot; huang2025interleaving introduces clear intermediate representations to decompose complex prompts into structured steps or explicit spatial layouts, improving compositional consistency and coherence. Another line of work reason2edit; mi2025thinkdiff; bagel2025 encourages models to reason about intent and constraints before drawing, moving beyond one-shot planning to better satisfy complex requirements, like R-Genie rgenie, which infers latent user intent instead of merely following the surface-level prompt. A third line of work guo2025canwegen; wu2025omnigen2; li2025reflectdit; li2025editthinker; yin2025reasonedit focuses on post-generation refinement by introducing reflection-and-correction mechanisms that assess the generated image, diagnose issues, and iteratively update the output to improve final quality. For example, Reflect-DiT li2025reflectdit introduces explicit self-reflection to guide revision, while EditThinker li2025editthinker enables reasoning via multi-round reflective interactions throughout the editing process. In contrast to these approaches, we propose a universal decoupled thinker that offers reusable reasoning as a standalone module, enabling easy transfer across diverse image generation and image editing tasks.
@@ 3 Data Construction
We aim to train a standalone Thinker that augments existing diffusion generators with transferable reasoning while remaining generator-agnostic. To this end, we construct HieraReason-40K, a selected general-purpose corpus by combining four sources that cover text-to-image generation, general image editing, reasoning image generation, and reasoning image editing tasks han2025unireditbench; huang2025interleaving; qian2025pico; fang2025fluxreason. Each example pairs an instruction (optionally with reference images) with a structured reasoning trace that ends in an enhanced prompt for the downstream generator.
@@
We use Gemini3-Pro nanobanana to generate initial structured reasoning traces, followed by automatic normalization to enforce strict format consistency (e.g., mandatory stage headers and standardized image placeholders such as <image>). We further filter or rewrite samples that violate the trace format, fail to follow the edit-only principle for image editing, produce non-visual or underspecified targets, or exhibit inconsistencies between the reasoning trace and the final enhanced prompt. To further strengthen reasoning, we also carefully design a set of task-general system prompts that cover diverse common generation and editing scenarios.
@@ 4.2 Dual-Phase Reinforcement Learning
While joint fine-tuning provides an initial alignment, it leaves a nontrivial reasoning–execution gap: the Thinker may produce plans that are plausible in text but suboptimal for the generator to execute. To address this without additional manual annotation, we introduce a dual-phase reinforcement learning strategy based on Group Relative Policy Optimization guo2025deepseekr1. The key idea is to sample multiple candidate traces for the same request, execute them with the generator, and train the Thinker by relative advantage feedback, promoting outputs that lead to better images and suppressing those that do not.
@@
Phase 2: Generation-Oriented RL. With the Thinker providing reliable plans, we next improve the Generator’s execution fidelity. However, probability-flow ODE sampling in diffusion models is essentially deterministic, limiting the stochastic rollouts required by reinforcement learning. Following a Flow-GRPO-like idea flowgrpo, we convert the ODE sampler into an equivalent reverse-time SDE to introduce controlled randomness, enabling $G$ distinct rollouts $\{z_{1},z_{2},\dots,z_{G}\}$ for the same instruction and optimizing the Generator accordingly:
@@ 5.1 Experimental Setup
Model configuration. Unless otherwise specified, Unified Thinker uses Qwen2.5-VL-7B, and we additionally report results with Qwen3-VL-8B bai2024qwenvl; bai2025qwen3vltechnicalreport.We use Qwen-Image-Edit qwenimage-edit as the base generator to execute the visual specifications produced by the Thinker. For reinforcement learning and automated evaluation, we adopt Qwen3-VL-30B bai2025qwen3vltechnicalreport as the reward model, which provides feedback on both visual correctness and logical consistency.
@@
Evaluation benchmarks. We evaluate on WiseBench wisebench, RISEBench risebench, GEditBench liu2025step1x-edit, and PRISMBench fang2025fluxreason. These benchmarks cover diverse knowledge domains and editing operations, requiring models to combine high-level semantic reasoning (e.g., temporal, and logical inference) with low-level visual manipulation (e.g., content preservation).
@@ 5.2 Main Results
Reasoning-based image editing (RISEBench). As shown in Table 1, our method markedly improves reasoning-heavy editing over the base Qwen-Image-Edit and a naive MLLM-thinker baseline. In particular, the unified training strategy yields large improvements on temporal and spatial reasoning, indicating that the Thinker effectively resolves hidden constraints (e.g., temporal shifts or relational edits) and reduces semantic drift during diffusion execution.
Text-to-image reasoning (WiseBench). Table 4 shows that Unified Thinker achieves the strongest overall performance among open-source models and improves most domain categories, substantially narrowing the gap to closed-source frontier models such as GPT-4o. Gains are especially notable in categories that demand precise entity grounding and knowledge retrieval (e.g., cultural and biology), suggesting that explicit planning helps translate implicit constraints into executable visual specifications.
General generation and editing quality. Beyond reasoning-centric benchmarks, we further confirm that incorporating the Thinker does not compromise general-purpose generation or editing performance. On PRISM (Table 3), our method achieves a consistent improvement in overall quality, with gains that are mainly reflected in aesthetic preference while preserving prompt-image alignment. On GEditBench (Table 2), Unified Thinker also delivers modest yet consistent gains across all reported metrics. Together, these results suggest that the planning stage improves instruction decomposition and visual target specification without weakening the Generator’s core rendering ability, and can even provide small benefits under standard, non-reasoning workloads.
@@ A.4 Reward Model and Design
We use a VLM-based reward model to provide a scalar supervision signal for both image-editing and text-to-image (T2I) training. For image editing, the judge is conditioned on the pre-edit image, the post-edit image, the edit instruction (edit_prompt), and a reference description of the intended outcome (edit_prompt_cot), and returns three integer subscores on a 1–5 scale: Appearance Consistency (whether non-instructed regions remain unchanged), Reasoning/Alignment (how well the edited image matches the intended result under the instruction), and Visual Plausibility (realism and overall generation quality); these subscores are aggregated into a single scalar reward. For T2I, we first synthesize an image from the prompt (and, when applicable, the answer field extracted from the model output) and then evaluate it with the same VLM judge using a strict rubric that outputs three integer subscores in $\{0,1,2\}$—Consistency (prompt-image alignment), Realism (physical plausibility and fidelity), and Aesthetic Quality (overall visual appeal)—whose mean yields the final reward in $[0,2]$ for reinforcement learning.
@@ A.5 Evaluated Comparative Models
Gemini-2.5-Flash-Image(Nano Banana): A state-of-the-art multimodal model by Google optimized for high-fidelity text-to-image generation, complex image editing, and multi-image composition nanobanana.
GPT-Image-1 & GPT-4o: OpenAI’s unified multimodal series that demonstrates advanced spatial-temporal reasoning and end-to-end processing across text and vision hurst2024gpt4o.
Gemini-2.0-Flash: A multimodal model from Google designed for real-time visual and textual reasoning tasks geminifamily.
BAGEL: A unified understanding and generation multimodal framework that incorporates Chain-of-Thought (CoT) reasoning to improve logical deduction in visual tasks bagel.
Qwen-Image / Edit: A series of vision-language models from Alibaba; the Edit variant is specifically fine-tuned for instruction-based image manipulation qwenimage-edit.
EMU2: A generative multimodal model that uses a unified modeling framework for both visual-sequential understanding and generation emu2.
FLUX.1 (Dev/Kontext): A flow-matching based rectified flow transformer model known for superior text rendering and adherence to complex prompts blackforest2024flux.
Stable Diffusion 3.5 (SD3.5): A Multimodal Diffusion Transformer (MMDiT) architecture optimized for high-resolution synthesis and prompt following sd3.5.
OmniGen: A unified image generation model capable of handling various tasks including generation, editing, and control within a single framework wu2025omnigen2.
Ovis: An open-source structural visual-language model designed to process high-resolution images with structural integrity ovis.
Step1X-Edit (v1.1/v1.2): A family of generative models by StepFun; the v1.2 variants utilize "thinking" and "reflection" mechanisms to improve reasoning-heavy editing tasks liu2025step1x-edit.
UniWorld (V1/V2): A multimodal world model framework designed for spatial-temporal understanding and high-fidelity video/image synthesis lin2025uniworld.
Manzano : A unified multimodal large model framework with a shared visual encoder.MANZANO.
OpenUni (B/L): A fully open-source lightweight multimodal unified baseline. It connects existing multimodal large language models with diffusion models through learnable queries and a lightweight Transformer connector, thereby enabling simultaneous multimodal understanding and image generation. openuni.
MetaQuery-XL: An expanded multimodal. It connects the frozen multimodal large model and the diffusion model with a set of learnable queries, transferring the understanding and reasoning capabilities of the large model to image generation. metaqueriesxl.
Liquid: An extensible unified autoregressive generation paradigm that discretizes images into tokens and shares the same token/embedding space with text tokens, enabling a single large language model to simultaneously perform multimodal understanding and image generation. liquid.

### v2

@@ 1 Introduction
The rapid evolution of diffusion-based foundation models Ho et al. (2020); Dhariwal and Nichol (2021); Rombach et al. (2022); Rafailov et al. (2023) has driven an unprecedented leap in high-fidelity image synthesis. Advanced proprietary models such as GPT-4o Hurst et al. (2024) and Nano Banana Comanici et al. (2025) have recently demonstrated strong reasoning-driven image generation under complex instructions. In contrast, despite steady progress in open-source systems Esser et al. (2024); Labs (2024); Deng et al. (2025); Wu et al. (2025a); Liao et al. (2025), current open-source models still exhibit a clear gap in handling logic-intensive or implicit directives Niu et al. (2025); Zhao et al. (2025); He et al. (2025); Liu et al. (2025b).
Current attempts to bridge this gap follow two primary approaches. Built-in Reasoning internalizes reasoning into the generator via unified training that couples multimodal understanding with generation Deng et al. (2025); Xie et al. (2025); Xiao et al. (2025). However, this tight entanglement reduces modularity and may destabilize training, often degrading the generator’s visual fidelity. In contrast, External Planner-Driven methods use an MLLM to plan for a mostly frozen generator Wu et al. (2025a); Lin et al. (2025); Li et al. (2025a); He et al. (2025). While modular, they suffer from a reasoning–execution mismatch: text-space plans are not grounded in the generator’s capabilities, so even correct plans can cause visual failures, and iterative planning further increases compute.
@@
However, this decoupled design still faces additional challenges: as shown in Fig. 1, without proper alignment, a naive Thinker may produce plausible reasoning that the Generator cannot execute. To bridge the reasoning-to-execution gap, we introduce a dedicated data-to-training pipeline to align planning with visual outcomes. We first construct HieraReason-40K, a hierarchical reasoning dataset synthesized with Gemini-3-Pro Comanici et al. (2025), which pairs complex instructions with structured, executable plans to teach the Thinker the desired planning format and basic logical decomposition. We then adopt a two-stage training strategy: we perform joint supervised fine-tuning on HieraReason-40K to establish initial plan quality, followed by an end-to-end dual-phase reinforcement learning procedure that places the Generator in the loop and optimizes the Thinker using rewards computed from the final image’s constraint satisfaction. This directly grounds the Thinker’s policy in pixel-level feedback, encouraging plans that are not only semantically plausible but also executable under the Generator’s capabilities.
@@ 2.1 Foundational Generative Models
Modern image generation is predominantly anchored in diffusion-based frameworks Ho et al. (2020); Rombach et al. (2022). Recent advances Esser et al. (2024); Labs (2024); Wu et al. (2025a) build upon Diffusion Transformers Peebles and Xie (2023) and flow matching Lipman et al. (2022) to improve fidelity, prompt alignment, and diversity in latent diffusion models. Meanwhile, an emerging direction unifies autoregressive modeling with visual generation in a single framework, giving rise to unified multimodal models Deng et al. (2025); Wu et al. (2025b); Xie et al. (2025); Xiao et al. (2025). For instance, Bagel Deng et al. (2025) uses a transformer backbone to jointly model text and image tokens, whereas OmniGen Xiao et al. (2025) dispenses with external encoders and handles multiple vision tasks through a unified pipeline. In parallel, image editing has evolved from mask-based inpainting Zhuang et al. (2024); Ju et al. (2024) to instruction-guided manipulation Brooks et al. (2023); Yu et al. (2025). To further enhance instruction following, recent methods Huang et al. (2024); Fu et al. (2024); Lin et al. (2025); Liu et al. (2025b) such as Qwen-Image-Edit Wu et al. (2025a) leverage MLLMs for instruction parsing and planning. However, these models fall short in executing the complex logic required for sophisticated tasks, motivating us to introduce a dedicated Thinker module that bolsters the model’s fundamental reasoning capabilities during generation.
@@ 2.2 Reasoning for Image Generation
Recent research has moved beyond the one-shot mapping paradigm by explicitly incorporating reasoning into the image generation process. One line of work Jiang et al. (2025); Wang et al. (2025); Liao et al. (2025); Zhou et al. (2026); Qin et al. (2025) introduces clear intermediate representations to decompose complex prompts into structured steps or explicit spatial layouts, improving compositional consistency and coherence. Another line of work He et al. (2025); Mi et al. (2025); Deng et al. (2025) encourages models to reason about intent and constraints before drawing, moving beyond one-shot planning to better satisfy complex requirements, like R-Genie Zhang et al. (2025), which infers latent user intent instead of merely following the surface-level prompt. A third line of work Guo et al. (2025b); Wu et al. (2025c); Li et al. (2025b, a); Yin et al. (2025) focuses on post-generation refinement by introducing reflection-and-correction mechanisms that assess the generated image, diagnose issues, and iteratively update the output to improve final quality. For example, Reflect-DiT Li et al. (2025b) introduces explicit self-reflection to guide revision, while EditThinker Li et al. (2025a) enables reasoning via multi-round reflective interactions throughout the editing process. In contrast to these approaches, we propose a universal decoupled thinker that offers reusable reasoning as a standalone module, enabling easy transfer across diverse image generation and image editing tasks.
@@ 3 Data Construction
We aim to train a standalone Thinker that augments existing diffusion generators with transferable reasoning while remaining generator-agnostic. To this end, we construct HieraReason-40K, a selected general-purpose corpus by combining four sources that cover text-to-image generation, general image editing, reasoning image generation, and reasoning image editing tasks Han et al. (2025); Huang et al. (2025); Qian et al. (2025); Fang et al. (2025). Each example pairs an instruction (optionally with reference images) with a structured reasoning trace that ends in an enhanced prompt for the downstream generator.
@@
We use Gemini3-Pro Comanici et al. (2025) to generate initial structured reasoning traces, followed by automatic normalization to enforce strict format consistency (e.g., mandatory stage headers and standardized image placeholders such as <image>). We further filter or rewrite samples that violate the trace format, fail to follow the edit-only principle for image editing, produce non-visual or underspecified targets, or exhibit inconsistencies between the reasoning trace and the final enhanced prompt. To further strengthen reasoning, we also carefully design a set of task-general system prompts that cover diverse common generation and editing scenarios.
@@ 4.2 Dual-Phase Reinforcement Learning
While joint fine-tuning provides an initial alignment, it leaves a nontrivial reasoning–execution gap: the Thinker may produce plans that are plausible in text but suboptimal for the generator to execute. To address this without additional manual annotation, we introduce a dual-phase reinforcement learning strategy based on Group Relative Policy Optimization Guo et al. (2025a). The key idea is to sample multiple candidate traces for the same request, execute them with the generator, and train the Thinker by relative advantage feedback, promoting outputs that lead to better images and suppressing those that do not.
@@
Phase 2: Generation-Oriented RL. With the Thinker providing reliable plans, we next improve the Generator’s execution fidelity. However, probability-flow ODE sampling in diffusion models is essentially deterministic, limiting the stochastic rollouts required by reinforcement learning. Following a Flow-GRPO-like idea Liu et al. (2025a), we convert the ODE sampler into an equivalent reverse-time SDE to introduce controlled randomness, enabling $G$ distinct rollouts $\{z_{1},z_{2},\dots,z_{G}\}$ for the same instruction and optimizing the Generator accordingly:
@@ 5.1 Experimental Setup
Model configuration. Unless otherwise specified, Unified Thinker uses Qwen2.5-VL-7B, and we additionally report results with Qwen3-VL-8B Bai et al. (2025b, a).We use Qwen-Image-Edit Wu et al. (2025a) as the base generator to execute the visual specifications produced by the Thinker. For reinforcement learning and automated evaluation, we adopt Qwen3-VL-30B Bai et al. (2025a) as the reward model, which provides feedback on both visual correctness and logical consistency.
@@
Evaluation benchmarks. We evaluate on WiseBench Niu et al. (2025), RISEBench Zhao et al. (2025), GEditBench Liu et al. (2025b), and PRISMBench Fang et al. (2025). These benchmarks cover diverse knowledge domains and editing operations, requiring models to combine high-level semantic reasoning (e.g., temporal, and logical inference) with low-level visual manipulation (e.g., content preservation).
@@ 5.2 Main Results
Reasoning image editing (RISEBench). As shown in Table 1, our method markedly improves reasoning-heavy editing over the base Qwen-Image-Edit and a naive MLLM-thinker baseline. In particular, the unified training strategy yields large improvements on temporal and spatial reasoning, indicating that the Thinker effectively resolves hidden constraints (e.g., temporal shifts or relational edits) and reduces semantic drift during diffusion execution.
Reasoning text-to-image (WiseBench). Table 4 shows that Unified Thinker achieves the strongest overall performance among open-source models and improves most domain categories, substantially narrowing the gap to closed-source frontier models such as GPT-4o. Gains are especially notable in categories that demand precise entity grounding and knowledge retrieval (e.g., cultural and biology), suggesting that explicit planning helps translate implicit constraints into executable visual specifications.
General generation and editing. Beyond reasoning-centric benchmarks, we further confirm that incorporating the Thinker does not compromise general-purpose generation or editing performance. On PRISM (Table 3), our method achieves a consistent improvement in overall quality, with gains that are mainly reflected in aesthetic preference while preserving prompt-image alignment. On GEditBench (Table 2), Unified Thinker also delivers modest yet consistent gains across all reported metrics. Together, these results suggest that the planning stage improves instruction decomposition and visual target specification without weakening the Generator’s core rendering ability, and can even provide small benefits under standard, non-reasoning workloads.
@@ A.4 Reward Model and Design
We use a VLM-based reward model to provide a scalar supervision signal for both image-editing and text-to-image (T2I) training. For image editing, the judge is conditioned on the pre-edit image, the post-edit image, the edit instruction (edit prompt), and a reference description of the intended outcome (reasoning edit prompt), and returns three integer subscores on a 1–5 scale: Appearance Consistency (whether non-instructed regions remain unchanged), Reasoning/Alignment (how well the edited image matches the intended result under the instruction), and Visual Plausibility (realism and overall generation quality); these subscores are aggregated into a single scalar reward. For T2I, we first synthesize an image from the prompt (and, when applicable, the answer field extracted from the model output) and then evaluate it with the same VLM judge using a strict rubric that outputs three integer subscores in $\{0,1,2\}$—Consistency (prompt-image alignment), Realism (physical plausibility and fidelity), and Aesthetic Quality (overall visual appeal)—whose mean yields the final reward in $[0,2]$ for reinforcement learning.
@@ A.5 Evaluated Comparative Models
Gemini-2.5-Flash-Image(Nano Banana): A state-of-the-art multimodal model by Google optimized for high-fidelity text-to-image generation, complex image editing, and multi-image composition Comanici et al. (2025).
GPT-Image-1 & GPT-4o: OpenAI’s unified multimodal series that demonstrates advanced spatial-temporal reasoning and end-to-end processing across text and vision Hurst et al. (2024).
Gemini-2.0-Flash: A multimodal model from Google designed for real-time visual and textual reasoning tasks Team et al. (2025).
BAGEL: A unified understanding and generation multimodal framework that incorporates Chain-of-Thought (CoT) reasoning to improve logical deduction in visual tasks Deng et al. (2025).
Qwen-Image / Edit: A series of vision-language models from Alibaba; the Edit variant is specifically fine-tuned for instruction-based image manipulation Wu et al. (2025a).
EMU2: A generative multimodal model that uses a unified modeling framework for both visual-sequential understanding and generation Sun et al. (2024).
FLUX.1 (Dev/Kontext): A flow-matching based rectified flow transformer model known for superior text rendering and adherence to complex prompts Labs (2024).
Stable Diffusion 3.5 (SD3.5): A Multimodal Diffusion Transformer (MMDiT) architecture optimized for high-resolution synthesis and prompt following Esser et al. (2024).
OmniGen: A unified image generation model capable of handling various tasks including generation, editing, and control within a single framework Wu et al. (2025c).
Ovis: An open-source structural visual-language model designed to process high-resolution images with structural integrity Lu et al. (2024).
Step1X-Edit (v1.1/v1.2): A family of generative models by StepFun; the v1.2 variants utilize "thinking" and "reflection" mechanisms to improve reasoning-heavy editing tasks Liu et al. (2025b).
UniWorld (V1/V2): A multimodal world model framework designed for spatial-temporal understanding and high-fidelity video/image synthesis Lin et al. (2025).
Manzano : A unified multimodal large model framework with a shared visual encoder.Li et al. (2025c).
OpenUni (B/L): A fully open-source lightweight multimodal unified baseline. It connects existing multimodal large language models with diffusion models through learnable queries and a lightweight Transformer connector, thereby enabling simultaneous multimodal understanding and image generation. Wu et al. (2025e).
MetaQuery-XL: An expanded multimodal. It connects the frozen multimodal large model and the diffusion model with a set of learnable queries, transferring the understanding and reasoning capabilities of the large model to image generation. Pan et al. (2025).
Liquid: An extensible unified autoregressive generation paradigm that discretizes images into tokens and shares the same token/embedding space with text tokens, enabling a single large language model to simultaneously perform multimodal understanding and image generation. Wu et al. (2025d).